Grade 11General Physics

two trains emerging towards each other with velocities 30km/hr are 60kms apart.A bird is sitting on a train which flies from the train to the other train with velocity 60km/hr. the bird continuosly flies from one train to the other . How many times will the bird fly here and there till the trains collide?

Rijit Singh
10 Years agoGrade 11

1 Answer

Saurabh Koranglekar
6 Years ago
Dear student

60/(30+30) = 1 hr

dis by bird = 1*60 = 60 km

Regards